OpenStudy (kimberly101):

Find the selling price for this item: A sound system costing $499 is sold at a loss of 45%.

OpenStudy (anonymous):

45% of 499= 0.45*499 =224.55 Selling price=Cost price-Loss

OpenStudy (anonymous):

so it's sold at 55% of the cost price then. find 55% of $499

OpenStudy (anonymous):

Use the formula: \[\large Selling \;Price = \frac{(100 - Loss \; Percentage)}{100} \times Cost \;Price\]

OpenStudy (kimberly101):

Thankyou

OpenStudy (anonymous):

\[S.P = \frac{(100-45)}{100} \times 499 = \frac{55}{100} \times 499\] Solve this..

OpenStudy (anonymous):

Welcome dear..